tritrations were performed on a 300.00mL samples of water taken from Lake Ontario. 4 trials were done. 54.5 mL/35.8 mL/ 53.8 mL/54.1 mL of 0.020 mol/L KOH was used to neutralize the sulphuric acid (H2SO4) in the lake water sample. Calculate the concentration of the phosphoric acid in the lake.


1
Expert's answer
2022-05-30T21:30:03-0400

H2SO4 + 2KOH = K2SO4 + 2H2O

1. We find the amount of substance (mol) of KOH used:

n = CM * V in the formula,

n — amount of substance (mol);

CM — concentration of molarity (mol/l) = =0,02 mol/l;

V — volume of solution (l) .

2. According to the above reaction, 1 mole of H2SO4 reacts with 2 moles of KOH. therefore we divide the amount of KOH (mol) by 2. resulting in an amount (mole) of H2SO4

3. Divide the result by the volume of the water sample being tested (0.3 l). resulting in a molar concentration of H2SO4

We calculate the above steps for each trials


a)

1. n = 0,02 * 0,0545 = 0,00109 mol KOH

2. 0,00109 / 2 = 0,000545 mol H2SO4

3. 0,000545 / 0,3 = 0,0018 mol/l

b)

1. n = 0,02 * 0,0358 = 0,000716mol KOH

2. 0,000716 / 2 = 0,000358 mol H2SO4

3. 0,000358 / 0,3 = 0,0012 mol/l

c)

1. n= 0,02 * 0,0538 = 0,001076 mol KOH

2. 0,0010076 / 2 = 0,000538 mol H2SO4

3. 0,000538 / 0,3 = 0,00179 mol/l

d)

1. n = 0,02 * 0,0541 = 0,001082 mol KOH

2. 0,001082 / 2 = 0,000541 mol H2SO4

3. 0,000541 / 0,3 = 0,0018 mol/l


To find of the concentration of H2SO4 in the lake, we find the average value of each trials result:

(0,0018+0,0012+0,00179+0,0018) / 4 = 0,0016475 mol/l

ANSWER: 0,0016475 mol/l H2SO4